# Agentic Structured Graph Traversal for Root Cause Analysis of Code-related Incidents in Cloud Applications

### v1

@@ Abstract
Cloud incidents pose major operational challenges in production, with unresolved production cloud incidents cost on average over \$2M per hour. Prior research identifies code- and configuration-related issues as the predominant category of root causes in cloud incidents. This paper introduces Praxis, an orchestrator that manages and deploys an agentic workflow for diagnosing code- and configuration-caused cloud incidents. Praxis employs an LLM-driven structured traversal over two types of graph: (1) a service dependency graph (SDG) that captures microservice-level dependencies; and (2) a hammock-block program dependence graph (PDG) that captures code-level dependencies for each microservice. Compared to state-of-the-art ReAct baselines, Praxis improves RCA accuracy by up to $3.1\times$ while reducing token consumption by $3.8\times$. Praxis is demonstrated on a set of 30 comprehensive real-world incidents that is being compiled into an RCA benchmark.
@@ I Introduction
Production cloud incidents cost on average over \$2M per hour [newrelic_2025_report, newrelic_2025]. While operational mitigation actions such as reboot, restart, scale-out, and rollback can resolve many such incidents, approximately 24% remained unresolved by operational mitigation [ghosh_2022_socc]. In those cases, engineers supported by tools must perform root-cause analysis (RCA)–a diagnostic process that identifies the origin (i.e., root causes) of the failure. When RCA is ineffective, outages can be prolonged and highly disruptive [monzo_2017_outage, reddit_pi_day_outage, aws_message_101925, prince_cloudflare_outage_2025].
This paper introduces Praxis, an orchestrator that manages and deploys an agentic workflow for diagnosing code- and configuration-caused cloud incidents. Praxis employs a large language model (LLM)-driven structured traversal over two types of graph: (1) a service dependency graph (SDG) that captures microservice-level dependencies; and (2) a hammock- block [johnson_pst_1994, fubo_hammock_2004] program dependence graph (PDG) that captures code-level dependencies for each microservice. Together, these graphs efficiently encode microservice- and code-level dependencies. By explicitly and jointly imposing both microservices and code dependency structures in the agentic RCA workflow, Praxis extends the diagnosis beyond observability symptoms, making it capable of identifying and explaining root causes that lie in code or configurations. We demonstrate the breadth of Praxis using 30 real-world examples, integrated into a comprehensive benchmark which is in the process of being open-sourced.
@@
    Praxis, an agentic approach for cloud incident RCA with structured, LLM-driven graph reasoning and traversal by utilizing microservices and program dependency graphs.
@@
    Code-Cloud-RCA Benchmark consisting of 30 different scenarios and each incorporating a unique software, configuration, deployment, or resource-related fault observed in the real-world [cbs_database, k8saf, monzo_2017_outage, reddit_pi_day_outage, ghosh_2022_socc, gunawi2014cloudbug, hotos_19], and injected in a live Kubernetes cloud environment.
@@
An important reason for Praxis’s effectiveness gain is its ability to reduce the context space and focus microservice- and code-level analysis on the regions most implicated by an incident, guided by the underlying dependency structure. Prior work [roy2024exploring_rca_agents, xu2025openrca, jha2025itbench, chen2025stratus, li2025coca_icse, wang2024rcagent] does not impose such a structure and relies entirely on the LLM’s built-in autoregressive analysis over unstructured, plain-text prompts, which is shown to be less effective. Praxis instead explicitly constrains the LLM’s reasoning along graph-defined dependencies encoded by PDG and SDG via LLM-driven graph traversal.
Praxis stands at the intersection of cloud incident diagnosis (an SRE task) and program analysis (an SWE task) [^1] SRE: site reliability engineering; SWE: software engineering.. By jointly investigating microservice-level and code-level dependencies–a process we refer to as Cross-SDG-PDG traversalPraxis can diagnose incidents whose fault chains interleave microservice and code dependencies across multiple hops. This unique capability distinguishes Praxis from existing approaches, which typically limit their investigation to either the microservice level using observability data [roy2024exploring_rca_agents, jha2025itbench, xu2025openrca, chen2025aiopslab, chen2025stratus, wang2024rcagent] or the code level based on engineer-curated postmortem issue tickets [li2025coca_icse]. We demonstrate this capability of Praxis in §III.
@@ II Terminology
Root cause analysis (RCA). Following the understanding espoused in [google_sre_2016], RCA is the diagnostic process that (a) identifies the microservice responsible for the incident, and (b) the statement(s), function(s), or configuration(s) identifying the origin of the incident, and (c) provide a concise reasoning of its manifestation to facilitate remediation.
Service dependency graph (SDG) Service dependency graph of a cloud application is a directed graph with nodes representing individual microservices or associated resources (e.g., pods, configMaps) and directed edges showing the dependency between them. In any given interval, the graph represents the current dynamics among the microservices.
Program dependence graph (PDG). A directed graph that captures microservice program structure [ferrante1987pdg], the nodes are hammock blocks [fubo_hammock_2004, johnson_pst_1994] and the edges are data, control, and call dependencies.
Hammock block. A hammock block is used to restructure unstructured code into a structured block format into a single-entry-single-exit (SESE) region [fubo_hammock_2004, johnson_pst_1994].
Observability. Observability with respect to microservices is the ability to understand the internal state of a distributed system by examining its generated data (e.g, logs, metrics, events, and distributed traces) via data analysis to gain real-time insights.
@@ III Approach Overview
This section introduces Praxis, an orchestrator that manages and deploys an agentic workflow for cloud incident RCA. At the core this approach is an LLM-driven structure traversal over two types of graph: (1) a SDG that captures microservice-level dependencies–providing a high-level localization of faulty microservices; and (2) a hammock-block PDG that captures code-level dependencies for each microservice–allowing for fine-grained RCA decisions that identify offending code paths and configurations responsible for the incident. Praxis’s hammock-block PDG provides an efficient structural representation of a microservice program, enabling seamless analysis of microservice code at different granularities. By integrating tools, observability data, and dependency graphs, Praxis delivers demonstratively accurate RCA for live cloud incidents, demonstrated on microservice-based Kubernetes applications. In summary the core functionalities of Praxis’s workflow include:
@@
Structure-aware agentic RCA. Upon triggering by active golden-signal alert(s), Praxis presents the relevant data (active alert(s) and error traces) to an LLM to identify and select microservices correlated with error signals in those data. The selected microservices become initial candidates for agentic focus. Subsequently, the LLM investigates each microservice by performing structure-aware reasoning through traversal of the corresponding hammock blocks. Praxis uniquely constructs the PDG using hammock blocks at different granularities spanning module, class, function, and statement levels, allowing the LLM to seamlessly move between levels as required to optimize analysis efficiency and accuracy. Leveraging this structure, the LLM iteratively traverses first the SDG and then the hammock-block PDG at the required level to narrow down the pool of potentially responsible microservices and the corresponding fault sites in code or configuration, until it reaches a conclusive decision on the root cause.
@@
Figure 1: Incident: Degraded external database returned empty responses, triggering a silent retry loop in the Recommendation service that manifested solely as high latency alert associated with the Recommendation service, without explicit error logs or error traces. Cross-SDG-PDG traversal: (1) LLM selects the Recommendation service for investigation based on the observed alert. (2) Investigation of the Recommendation service reveals a silent retry loop with missing error logs. Code traversal shows that an unresponsive External Product Database triggers this loop. (3) Subsequent investigation shows a storage failure in the External Product DB as the root cause. (4) The ReAct agent baseline fails to pinpoint the precise root cause, whereas Praxis successfully isolates the root cause.
Cross-SDG-PDG traversal: a challenging scenario. A key innovation of our approach is the ability to handle incidents where multiple microservices are identified as potential root causes. In such cases, the LLM must traverse from the first microservice into its specific hammock block within the PDG, and then transition to a second microservice according to the SDG and its respective hammock blocks to form a coherent and holistic RCA decision. Praxis facilitates this by attaching the PDG of each microservice to the global SDG, ensuring the LLM can perform these cross-layer traversals when needed. We demonstrate the necessity of this unique capability in Figure˜1, using a challenging incident that the baseline ReAct agent failed to solve.
As shown by the agent output (Figure˜1- 4), the ReAct baseline (equipped with SRE and SWE tools) concludes its diagnosis at the Recommendation service after inspecting its code (Figure˜1- 2). In contrast, Praxis executes a cross-SDG-PDG traversal: it pivots its investigation from code-level (PDG) back to the microservice-level (SDG) to identify the external database dependency (Figure˜1- 3), where it repeats the investigation process to identify the failed database as the root cause, aligning with the groundtruth. This example demonstrates that merely equipping a ReAct SRE agent with SWE/code tools, or treating the two domains independently (e.g., by straightforwardly composing SRE and SWE agents), yields suboptimal results. Those approaches fail because they lack a structure in decision-making to explicitly enforce the investigation of the next dependencies, causing them to prematurely conclude the investigation before arriving at the actual root cause–a critical gap that Praxis addresses.
@@ IV Praxis: Methodology
Phase 1 (§IV-AFigure˜2): Data gathering and dependence graph construction. Praxis collects the incident context: error traces from Jaeger [jaeger_2025], sustained alerts from Prometheus [prometheus_2025], the SDG snapshot from the cloud topology monitoring platform [datadog_2025], and PDG from the microservice source code using Tree-sitter [tree_sitter_zenodo_2025] and CLDK [rahul2024cldk].
@@ IV-A Phase 1: Data Gathering
    Distributed error traces. Distributed error traces record runtime request latencies and propagations of inter-service requests, explicitly highlighting error dependency paths across microservice boundaries.
@@
    Logs. Cloud application and microservices log entries and their associated error/warning patterns [^2] To avoid context overflow, only the $K$ most recent entries are collected. We set $K=40$ for our evaluation, as empirical testing showed that increasing the value further does not significantly alter RCA accuracy.. These logs provide rich runtime context (e.g., information, warnings, errors, stack traces, exceptions), offering direct insight into the timing and nature of runtime executions, errors, and failures.
@@
    Events. Kubernetes events regarding deployment updates and lifecycle changes that affect the microservice or its resources (e.g., ReplicaSet scaling, container restarts), providing diagnostic context on deployment/resource configuration changes that could reveal the root cause.
@@
    Metrics. Aggregated error and warning counts derived from logs and events, alongside span and trace latency measurements to quantify the severity of observed anomalies.
Praxis aggregates sustained alerts and error traces into the incident context, which serves as the initial description of the cloud incident. While pre-collected, logs, metrics, and events will be used in Phase 3 for per-microservice investigation.
@@
To perform fine-grained diagnosis, Praxis augments each microservice node in the SDG with a program dependence graph (PDG) [ferrante1987pdg], enabling reasoning over code paths and potential fault sites. For a microservice program $P$ comprising a finite set of single-entry–single-exit (SESE) hammock blocks $B=\{b_{1},\ldots,b_{n}\}$ [johnson_pst_1994, fubo_hammock_2004], the PDG is defined as a labeled directed graph $G_{P}=(V,E,\lambda_{V},\lambda_{E},\prec)$. Here, $V=B$, and edges $E=E_{\text{ctl}}\cup E_{\text{data}}\cup E_{\text{call}}$ capture control-, data-, and call-dependence between blocks; $\lambda_{V}$ and $\lambda_{E}$ annotate nodes and edges with their corresponding semantic attributes. A containment relation $\prec$ specifies when one hammock block syntactically contains another, inducing a hierarchical structure spanning module-, class-, function-, and statement-level regions. This hierarchy allows Praxis to traverse and reason about the program at whatever granularity is most informative for RCA.
@@
Figure˜2 shows a representative PDG of the Recommendation service from [opentelemetry_demo], comprising function- and statement-level hammock blocks. [^4] We omit the full PDG for brevity, as the actual structure contains 44 hammock blocks. Exp denotes expression statements (a block of contiguous, straight-line statements without branching), whereas other nodes represent branching logic (e.g., if/else). Function-level blocks are explicitly labeled with function names (e.g., ListRecommendation()).
@@ IV-B Phase 2: Microservice Candidate(s) Selection
In Phase 2 (Figure˜3), Praxis invokes an LLM to select an initial set of cloud microservice candidates potentially responsible for the incident to start the investigation process using the SDG and the incident context from Phase 1. Praxis prompts the LLM (denoted as $\mathcal{M}$) with $\psi_{{select}}$, conditioning the selection on the SDG $G_{C}=\left(V,E\right)$ and the incident context, which comprises active sustained alert(s) ($A$) and collected error traces ($T$). The LLM suggests a set of initial microservice candidates (up to NN [^5] NN is a user-defined variable that we set to $N=5$ for our experiments. Further increases of this value yielded no improvements to our results.) from nodes in $G_{C}$, and outputs an initial investigation narrative $I_{0}$ describing the observed alerts, traces, and the reasons why they were selected. In addition, Praxis initializes an investigation queue $Q$ to actively keep track of microservices that need to be investigated and initializes $Q$ with those selected candidates: $Q\leftarrow\mathcal{M}(\psi_{{select}}(A,T,\lambda_{V}(G_{C})))$.
@@
Figure 4: Praxis Phase 3: RCA decision-making. This process is repeated for the next focal entity that is (a) a dependee of the current focal entity and/or (b) suggested by the LLM based on the focal entity’s RCA decision.

### v2

# Praxis: Integrating Program Analysis with Observability for Root-Cause Analysis

@@ Abstract
Unresolved production cloud incidents cost an average of over \$2M per hour. This paper introduces Praxis, an orchestrator that manages and deploys an agentic workflow for diagnosing code- and configuration-caused cloud incidents. Praxis employs an LLM-driven structured traversal over two types of graph: (1) a service dependency graph (SDG) that captures microservice-level dependencies; and (2) a hammock-block program dependence graph (PDG) that captures code-level dependencies for each microservice. Compared to state-of-the-art ReAct baselines, Praxis improves RCA accuracy by up to $6.3\times$ while reducing token consumption by $5.3\times$. Praxis is demonstrated on a set of 30 comprehensive real-world incidents that is being compiled into an RCA benchmark.
publicationid: pubid:   © 2026 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. Accepted to appear in The 56th Annual IEEE/IFIP International Conference on Dependable Systems and Networks.
@@ I Introduction
Production cloud incidents cost an average of over \$2M per hour [newrelic_2025_report, newrelic_2025]. While operational mitigation actions such as reboot, restart, scale-out, and rollback can resolve many such incidents, approximately 24% are not resolved by operational mitigation [ghosh_2022_socc]. In those cases, engineers supported by tools must perform root-cause analysis (RCA), a diagnostic process that identifies the origin (i.e., root causes) of a failure. When RCA is ineffective, outages can be prolonged and highly disruptive [monzo_2017_outage, reddit_pi_day_outage, aws_message_101925, prince_cloudflare_outage_2025].
This paper introduces Praxis, an orchestrator that manages and deploys an agentic workflow for diagnosing code- and configuration-caused cloud incidents. Praxis employs an LLM-driven structured traversal over two types of graph: (1) a service dependency graph (SDG) that captures microservice-level dependencies; and (2) a hammock- block [johnson_pst_1994, fubo_hammock_2004] program dependence graph (PDG) that captures code-level dependencies for each microservice. Together, these graphs efficiently encode microservice- and code-level dependencies. By explicitly and jointly imposing both microservices and code dependency structures in the agentic RCA workflow, Praxis extends the diagnosis beyond observability symptoms, making it capable of identifying and explaining root causes that lie in code or configurations. We demonstrate the breadth of Praxis using 30 real-world examples, integrated into a comprehensive benchmark that is open-sourced.
@@
    Praxis, an agentic approach for cloud incident RCA with structured, LLM-driven graph reasoning and traversal that utilizes microservices and program dependency graphs.
@@
    Code-Cloud-RCA Benchmark, integrated into ITBench [itbench_k8s_topology_monitor_2025], consisting of 30 different scenarios, each incorporating a unique software, configuration, deployment, or resource-related fault observed in the real world [cbs_database, k8saf, monzo_2017_outage, reddit_pi_day_outage, ghosh_2022_socc, gunawi2014cloudbug, hotos_19] and injected into a live Kubernetes cloud environment.
@@
An important reason for Praxis’s effectiveness gain is its ability to reduce the context space and focus microservice- and code-level analysis on the regions most implicated by an incident, guided by the underlying dependency structure. Prior work [roy2024exploring_rca_agents, xu2025openrca, jha2025itbench, chen2025stratus, wang2024rcagent, li2025coca_icse] does not impose such a structure and relies entirely on the LLM’s built-in autoregressive analysis over unstructured, plain-text prompts, which is shown to be less effective. Praxis instead explicitly constrains the LLM’s reasoning along graph-defined dependencies via LLM-driven graph traversal.
Praxis implementation and benchmark scenarios used in this paper are publicly available at https://doi.org/10.5281/zenodo.19163486.
@@ II Terminology
Root cause analysis (RCA). Following the understanding espoused in [google_sre_2016], RCA is the diagnostic process that (1) identifies the microservice responsible for the incident and (2) the statement(s), function(s), or configuration(s) identifying the origin of the incident, and (3) provides concise reasoning about its propagation and manifestation to facilitate remediation.
Service dependency graph (SDG). The service dependency graph of a cloud application is a directed graph with nodes representing individual microservices and directed edges showing the dependency between them. In any given interval, the graph represents the current dynamics among the microservices.
Program dependence graph (PDG). A directed graph that captures microservice program structure [ferrante1987pdg]; the nodes are hammock blocks [fubo_hammock_2004, johnson_pst_1994] and the edges are data, control, and call dependencies.
Hammock block. A hammock block restructures unstructured code into a structured block corresponding to a single-entry, single-exit (SESE) region in code [fubo_hammock_2004, johnson_pst_1994].
Observability. In microservices, observability is the ability to infer the distributed system’s internal state from its generated data (e.g., logs, metrics, events, and distributed traces), enabling real-time insight through analysis.
@@ III Approach Overview
This section introduces Praxis, an orchestrator that manages and deploys an agentic workflow for cloud incident RCA. At the core of this approach is an LLM-driven structure traversal over two types of graph: (1) an SDG that captures microservice-level dependencies (providing a high-level localization of faulty microservices); and (2) a hammock-block PDG that captures code-level dependencies for each microservice (allowing for fine-grained RCA decisions that identify offending code paths and configurations responsible for the incident). Praxis’s hammock-block PDG provides an efficient structural representation of a microservice program, enabling seamless analysis of microservice code at different granularities. By integrating tools, observability data, and dependency graphs, Praxis delivers demonstratively accurate RCA for live cloud incidents, demonstrated on microservice-based Kubernetes applications. In summary, the core functionalities of Praxis’s workflow include the following.
@@
Structure-aware agentic RCA. After being triggered by active golden-signal alert(s), Praxis presents the relevant data (active alert(s) and error traces) to an LLM to identify and select microservices correlated with error signals in those data. The selected microservices become initial candidates for agentic focus. Subsequently, the LLM investigates each microservice by performing structure-aware reasoning through traversal of the corresponding hammock blocks. Praxis uniquely constructs the PDG using hammock blocks at different granularities spanning module, class, function, and statement levels, allowing the LLM to seamlessly move between levels as required to optimize analysis efficiency and accuracy. Leveraging this structure, the LLM iteratively traverses first the SDG and then the hammock-block PDG at the required level to narrow down the pool of potentially responsible microservices and the corresponding fault sites in code or configuration, until it reaches a conclusive decision on the root cause.
@@
Figure 1: Incident: Degraded external database returned empty responses, triggering a silent retry loop in the Recommendation service that manifested solely as a high-latency alert associated with the Recommendation service, without explicit error logs or error traces. Cross-SDG-PDG traversal: (1) LLM selects the Recommendation service for investigation based on the observed alert. (2) Investigation of the Recommendation service reveals a silent retry loop with missing error logs. Code traversal shows that an unresponsive External Product Database triggers this loop. (3) Subsequent investigation shows a storage failure in the External Product DB as the root cause. (4) The ReAct agent baseline fails to pinpoint the precise root cause, whereas Praxis successfully isolates the root cause.
Cross-SDG-PDG traversal: a challenging scenario. A key innovation of our approach is the ability to handle incidents for which multiple microservices are identified as potential root causes. In such cases, the LLM must traverse from the first microservice into its specific hammock block within the PDG, and then transition to a second microservice, according to the SDG and its respective hammock blocks, to form a coherent and holistic RCA decision. Praxis facilitates this by attaching the PDG of each microservice to the global SDG, ensuring that the LLM can perform these cross-layer traversals when needed. We demonstrate the necessity of this unique capability in Figure˜1, using a challenging incident that the baseline ReAct agent failed to solve.
@@ IV Praxis: Methodology
Phase 1 (§IV-AFigure˜2): Data gathering and dependence graph construction. Praxis collects the incident context: error traces from Jaeger [jaeger_2025], sustained alerts from Prometheus [prometheus_2025], the SDG snapshot from the cloud topology monitoring platform [datadog_2025], and the PDG, constructed using Tree-sitter [tree_sitter_zenodo_2025] and CLDK [rahul2024cldk], from the microservice source code.
@@ IV-A Phase 1: Data Gathering
    Distributed error traces. Distributed error traces, obtained via cloud-native observability tools (e.g., OpenTelemetry and Jaeger), record runtime request latencies and propagations of inter-service requests, explicitly highlighting error dependency paths across microservice boundaries for initial fault localization.
@@
    Logs. Cloud application and microservices log entries and their associated error/warning patterns [^1] To avoid context overflow, only the $K$ most recent entries are collected. We set $K=40$ for our evaluation, as empirical testing showed that increasing the value further does not significantly alter RCA accuracy. provide rich runtime context (e.g., information, warnings, errors, stack traces, exceptions), offering direct insight into the timing and nature of runtime executions, errors, and failures.
@@
    Events. Kubernetes events regarding deployment updates and lifecycle changes that affect the microservice or its resources (e.g., ReplicaSet scaling, container restarts) provide diagnostic context on deployment/resource configuration changes that could reveal the root cause.
@@
    Metrics include aggregated error and warning counts derived from logs and events, alongside span and trace latency measurements to quantify the severity of observed anomalies.
Praxis aggregates sustained alerts and error traces into the incident context, which serves as the initial description of the cloud incident. Moreover, logs, metrics, and events will be used in Phase 3 for per-microservice investigation.
@@
To perform fine-grained diagnosis, Praxis augments each microservice node in the SDG with a program dependence graph (PDG) [ferrante1987pdg], enabling reasoning over code paths and potential fault sites. For a microservice program $P$ comprising a finite set of single-entry–single-exit (SESE) hammock blocks $B=\{b_{1},\ldots,b_{n}\}$ [johnson_pst_1994, fubo_hammock_2004], the PDG is defined as a labeled directed graph $G_{P}=(V,E,\lambda_{V},\lambda_{E},\prec)$. Here, $V=B$, and edges $E=E_{\text{ctl}}\cup E_{\text{data}}\cup E_{\text{call}}$ capture control-, data-, and call-dependence between blocks. $\lambda_{V}$ and $\lambda_{E}$ annotate nodes and edges with their corresponding semantic attributes: $\lambda_{V}$ includes node type, defined and used variables, string literals, and code snippets, and $\lambda_{E}$ includes dependency type (control-flow, data-flow, or caller-callee) and associated variable or function names. A containment relation $\prec$ specifies when one hammock block syntactically contains another, inducing a hierarchical structure over module-, class-, function-, and statement-level regions. This hierarchy allows Praxis to traverse and reason about the program at the most informative granularity for RCA.
@@
Figure˜2 shows a representative PDG of the Recommendation service from [opentelemetry_demo], comprising function- and statement-level hammock blocks. [^3] We omit the full PDG for brevity, as the actual structure contains 44 hammock blocks. Exp denotes expression statements (blocks of contiguous, straight-line statements without branching), whereas other nodes represent branching logic (e.g., the try-catch node). Function-level blocks are explicitly labeled with function names (e.g., the GetProdList() node).
@@ IV-B Phase 2: Microservice Candidate(s) Selection
In Phase 2 (Figure˜3), Praxis invokes an LLM to select an initial set of cloud microservice candidates potentially responsible for the incident to start the investigation process using the SDG and the incident context from Phase 1. Praxis prompts the LLM (denoted by $\mathcal{M}$) with $\psi_{{select}}$, conditioning the selection on the SDG $G_{C}=\left(V,E\right)$ and the incident context, which comprises active sustained alert(s) ($A$) and collected error traces ($T$). The LLM suggests a set of initial microservice candidates (up to NN [^4] NN is a user-defined variable that we set to $N=5$ for our experiments. Further increases of this value yielded no improvements to our results.) from nodes in $G_{C}$, and outputs an initial investigation narrative $I_{0}$ describing the observed alerts, traces, and the reasons why they were selected. In addition, Praxis initializes an investigation queue $Q$ to actively keep track of microservices that need to be investigated and initializes $Q$ with those selected candidates: $Q\leftarrow\mathcal{M}(\psi_{{select}}(A,T,\lambda_{V}(G_{C})))$.
@@
Figure 4: Praxis Phase 3: RCA decision-making. This process is repeated for the next focal entity that is (1) a dependee of the current focal entity and/or (2) suggested by the LLM based on the focal entity’s RCA decision.